I m not able to solve this question. Can u please answer with detailed solution:The density of a mixture of O2 and N2 at STP is 1.3g/litre. Calculate the partial pressure of O2.

Grade:11

2 Answers

Pratham Ashish
14 years ago

hi,

let the total no. of moles be n in which n1 are of O2 & n2 are of N2

so,   n= n1 + n2

 let the volume be V ,

so density = total mass/ V

       1.3   = (n1 *32  + n2 *28 )/V  ...................1

at stp , Volume of one mole gas is 22.4 lt.

so the volume of n= n1 +n2 moles =  22.4 n

                                               V  = 22.4 (n1 + n2 )    putting in eq 1

     1.3   = (n1 *32  + n2 *28 )/ 22.4 (n1 + n2 ) 

    29.12 n1 + 29.12 n2 =  32 n1 + 28 n2

                       1.12 n2 = 2.88 n1

         n1/n2 = .389

 so, n1/n = .2801                     ( n2/n1 + 1 = (n1 +n2) /n1)

at stp total P = 101.325 k Pa

partial pressure of O2 = n1/n* P

                               = .2801 *101.325

                               = 28.37 k Pa

 

 

Nikit Saraf
14 years ago

Thank You!! so much for the reply..